OpenStudy (anonymous):

Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. the 2nd equation is –2x + 3y = -6 and if i wanted to put it in slope intercept form wouldnt it be (y=mx+b) so 3y=-2x+-6 or is that wrong and wouldnt the slope be -2 and the y intercept form be -6 ? correct me if im wrong thank you

OpenStudy (hba):

y=mx+c*

OpenStudy (anonymous):

-2x + 3y = -6 (add 2x to both sides) 3y = 2x - 6 (now divide both sides by 3) y = 2/3x - 2 <-- slope intercept form y = mx + b is slope intercept form. The slope will be in the m position and the y intercept will be in the b position. Therefore, the slope = 2/3 and the y intercept = -2

OpenStudy (anonymous):

do you have any questions because I will be more then happy to explain it

OpenStudy (anonymous):

actually no you explained it all thank so much

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)